### v2

@@ 3 The space of absolutely continuous sections
for each $i\in\{1,...,n\}$. We endow ${\Gamma_{AC}(\eta)}$ with the Hausdorff locally convex vector topology which is generated by the initial topology with respect to the mappings
$h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto h_{i}(\sigma)=d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}$
@@
$\Phi_{\eta,P}:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto\left(d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
@@
We will show that the image set of $\Phi_{\eta,P}$, denoted by $\text{Im}(\Phi_{\eta,P})$, coincide with the closed subspace $K$. Indeed, the space $\text{Im}(\Phi_{\eta,P})$ is contained in $K$ by definition of $\Phi_{\eta,P}$. Let us consider now $\tau=(\tau_{i})_{i=1}^{n}\in K$. We define the maps
@@
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. If $E$ is a Banach space (resp. a Fréchet space), then ${\Gamma_{AC}(\eta)}$ is a Banach space (resp. a Fréchet space).
@@
This follows from the fact that each vector space $AC_{L^{p}}([t_{i-1},i],E)$ is a Banach space (resp. a Fréchet space). ∎
@@
Let $E_{M}$ and $E_{N}$ be the modeled spaces of $M$ and $N$ respectively. By Lemma 2.15 we have $f\circ\eta\in AC_{L^{p}}([a,b],N)$ and $Tf\circ\sigma\in AC_{L^{p}}([a,b],TN)$ for each $\sigma\in\Gamma_{AC}(\eta)$. Since $T_{\eta(t)}f\left(\sigma(t)\right)\in T_{f\circ\eta(t)}N$ for each $t\in[a,b]$, we have
@@
Thus $\widetilde{f}(\sigma)\in\Gamma_{AC}(f\circ\eta)$ for each $\sigma\in{\Gamma_{AC}(\eta)}$. The linearity of $\widetilde{f}$ is trivial.
Without loss of generality, we can chose a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ such that there is exists the families of charts $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ and $\{(\phi_{i},V_{i}):i\in\{1,..,n\}\}$ that verify the definition of absolutely continuous for $\eta$ and $f\circ\eta$ respectively, such that
@@
For $\sigma\in{\Gamma_{AC}(\eta)}$ and for each $i\in\{1,...,n\}$ we denote $\sigma_{i}=\sigma|_{[t_{i-1},t_{i}]}$ and we define the maps
$F_{i}:AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\tau\mapsto\left(d\phi_{i}\circ Tf\circ d\varphi_{i}^{-1}\right)\circ\tau$
@@
$F:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\tau\mapsto\left(\left(d\phi_{i}\circ Tf\circ d\varphi_{i}^{-1}\right)\circ\tau\right)_{i=1}^{n}$
@@
Let $i\in\{1,..,n-1\}$, if
$\tau_{i}:=F_{i}(d\varphi_{i}\circ\sigma_{i})=d\phi_{i}\circ Tf\circ\sigma_{i}$
@@
$d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),\tau_{i+1}(t_{i})\Big{)}$ $=d\varphi_{i}\circ(T\varphi_{i+1})^{-1}\Big{(}\varphi_{i+1}\circ\eta(t_{i}),d\phi_{i+1}\circ Tf\circ\sigma_{i+1}(t_{i})\Big{)}$
$=d\varphi_{i}\Big{(}\eta(t_{i}),Tf\circ\sigma_{i+1}(t_{i})\Big{)}$
$=d\varphi_{i}\circ Tf\circ\sigma_{i+1}(t_{i})$
$=d\varphi_{i}\circ Tf\circ\sigma_{i}(t_{i})$
@@
The topology of ${\Gamma_{AC}(\eta)}$ does not depend on the partition or charts chosen. Indeed, since the identity map $\text{id}_{M}:M\to M$ is smooth, by previous proposition the map
@@
is smooth regardless of the partition or charts chosen. Moreover, this map coincide with the identity map $\text{id}_{\Gamma}:{\Gamma_{AC}(\eta)}\to{\Gamma_{AC}(\eta)}$, $\sigma\mapsto\sigma$.
@@
are continuous [24, Lemma 3.2], the inclusion map $J_{\Gamma}:\Gamma_{AC}(\gamma)\to\Gamma_{C}(\gamma)$ is also continuous.
This implies that set
@@
Let $N_{1}$ and $N_{2}$ be smooth manifolds modeled on sequentially complete locally convex spaces, $p\in[1,\infty]$ and $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the $i$-projection for $i\in\{1,2\}$. If $\eta_{1}\in AC_{L^{p}}([a,b],N_{1})$ and $\eta_{2}\in AC_{L^{p}}([a,b],N_{2})$ then the map
$\widetilde{(\operatorname{pr}_{1},\operatorname{pr}_{2})}:\Gamma_{AC}(\eta_{1}\times\eta_{2})\to\Gamma_{AC}(\eta_{1})\times\Gamma_{AC}(\eta_{1}),\quad\sigma\mapsto(T\operatorname{pr}_{1},T\operatorname{pr}_{2})(\sigma)$
@@
This is consequence of Proposition 3.5, with $f=(\operatorname{pr}_{1},\operatorname{pr}_{2})$ and considering that the map $Tf=(T\operatorname{pr}_{1},T\operatorname{pr}_{2}):T(N_{1}\times N_{2})\to TN_{1}\times TN_{2}$ is a smooth diffeomorphism. ∎
@@
$\rho:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}\Gamma_{AC}(\eta|_{[t_{i-1},t_{i}]}),\quad\sigma\mapsto(\sigma|_{[t_{i-1},t_{i}]})_{i=1}^{n}$
@@
Let $j\in\{1,...,n\}$, If $P_{j}=\{t_{i-1},t_{i}\}$, since $\eta_{j}=\eta|_{[t_{j-1},t_{j}]}\in AC_{L^{p}}([t_{j-1},t_{j}],N)$, we have
$\rho=\Big{(}\Phi_{\eta_{i},P_{i}}^{-1}\Big{)}_{i=1}^{n}\circ\Phi_{\eta,P}.$
@@
$(\rho|^{\text{Im}(\rho)})^{-1}=\Phi_{\eta,P}^{-1}\circ\Big{(}\Phi_{\eta_{i},P_{i}}\Big{)}_{i=1}^{n}.$
@@
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. If $g:[c,d]\to[a,b]$ a the map as in Lemma 2.14, then $\eta\circ g\in AC_{L^{p}}([c,d],N)$ for each $\eta\in AC_{L^{p}}([a,b],N)$. Moreover, if $\eta\in AC_{L^{p}}([a,b],N)$, then the map
@@
Let $\eta\in AC_{L^{p}}([a,b],N)$, we will show first that $\eta\circ g\in AC_{L^{p}}([c,d],N)$. Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $N$ that verify the definition of absolute continuity for $\eta$. Since $g$ is a strictly increasing function, we can define a partition $Q=\{s_{1},...,s_{n}\}$ of $[c,d]$ such that $g(s_{i})=t_{i}$, for each $i\in\{1,...,n\}$. Moreover, given that
@@
which are continuous by Lemma 2.14. Considering the topological embeddings $\Phi_{\eta,P}$ and $\Phi_{\eta\circ g,Q}$ (as Proposition 3.2) with the same family of charts, if $(\tau_{i})_{i=1}^{n}\in\text{Im}(\Phi_{\eta,P})$ we have
@@
for each $i\in\{1,...,n-1\}$. If $G=(G_{1},...,G_{n})$, then $\left(G\circ\Phi_{\eta,P}\right)(\sigma)\in\text{Im}(\Phi_{\eta\circ g,Q})$ and
@@
Hence $L_{g}$ is continious and clearly linear. ∎
@@
is continuous and linear in the first argument. Moreover, for each $t\in[a,b]$ the point evaluation map
@@
is continuous and the evaluation map $\tilde{\epsilon}_{t}:\Gamma_{C}(\eta)\to TN$, $\sigma\mapsto\sigma(t)$ is smooth for each $t\in[a,b]$ (see [2]). Then $\epsilon=\tilde{\epsilon}\circ(J_{\Gamma},\text{Id}_{\mathbb{R}})$ and $\epsilon_{t}=\tilde{\epsilon}_{t}\circ J_{\Gamma}$, where $J_{\Gamma}:{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$ is the inclusion map, which is continuous linear by Remark 3.7. ∎
@@ 4 Manifolds of absolutely continuous functions
If a smooth manifold $N$ admits a local addition, then also its tangent manifold $TN$ admits a local addition [2, Lemma A.11]. Moreover, each manifold which admits local addition also admits a normalized local addition [2, Lemma A.14]. From now we will assume that each local addition is normalized.
@@
$\mathcal{V}_{\eta}:=\{\sigma\in{\Gamma_{AC}(\eta)}:\sigma([a,b])\subseteq\Omega\}.$
@@
The following Lemma is just an application of [3, Lemma 10.1] to our particular case.
@@
Let $E$ and $F$ sequentially complete locally convex spaces, $U\subseteq E$ open and $f:U\to F$ a map. If $F_{0}\subseteq F$ is a closed subset and $f(U)\subseteq F_{0}$, then $f:U\to F$ is smooth if and only if $f|^{F_{0}}:U\to F_{0}$ is smooth.
@@
We endow $AC_{L^{p}}([a,b],N)$ with the final topology with respect to the family of maps $\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta}$, for each $\eta\in AC_{L^{p}}([a,b],N)$. If we define the maps $\Psi_{\eta}^{C}\colon{\mathcal{V}}_{\eta}^{C}\to{\mathcal{U}}_{\eta}^{C}$ on the space of continuous functions $C([a,b],N)$ for each $\eta\in C([a,b],N)$ as Remark 4.3, then the final topology on $C([a,b],N)$ coincide with its topology. Hence the inclusion map
@@
$\mathcal{U}_{J(\eta)}^{C}:=\{\gamma\in C([a,b],N):(J(\eta),\gamma)([a,b])\subseteq\Omega^{\prime}\}$
is open in $C([a,b],N)$, thus

### v4

@@ 2 Preliminaries
and for each $i\in\{1,...,n\}$. We endow ${\Gamma_{AC}(\eta)}$ with the Hausdorff locally convex vector topology which is the initial topology with respect to the linear mappings
$h_{i}:{\Gamma_{AC}(\eta)}\to AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto h_{i}(\sigma)=d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}$ (2.5)
@@
$\Phi_{\eta,P}:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E),\quad\sigma\mapsto\left(d\varphi_{i}\circ\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$ (2.6)
@@
We will show that the image of $\Phi_{\eta,P}$, denoted by $\text{Im}(\Phi_{\eta,P})$, coincides with the closed subspace $K$. Indeed, the space $\text{Im}(\Phi_{\eta,P})$ is contained in $K$ by definition of $\Phi_{\eta,P}$. Let us consider now $\tau=(\tau_{i})_{i=1}^{n}\in K$. We define the maps
@@
Let $N$ be a smooth manifold modeled on a $E$ Banach space (resp. Frechet space), $p\in[1,\infty]$ and $\eta\in AC_{L^{p}}([a,b],N)$. Then the vector space ${\Gamma_{AC}(\eta)}$ is a Banach space (resp. Frechet space).
@@
This follows from the fact that each vector space $AC_{L^{p}}([t_{i-1},i],E)$ is a Banach space (resp. Frechet space) ∎
@@
Let $E_{M}$ and $E_{N}$ be the modeling spaces of $M$ and $N$ respectively. By Lemma 2.16 we have $f\circ\eta\in AC_{L^{p}}([a,b],N)$ and $Tf\circ\sigma\in AC_{L^{p}}([a,b],TN)$ for each $\sigma\in\Gamma_{AC}(\eta)$. Since $T_{\eta(t)}f\left(\sigma(t)\right)\in T_{f\circ\eta(t)}N$ for each $t\in[a,b]$, we have
@@
Thus $\widetilde{f}(\sigma)\in\Gamma_{AC}(f\circ\eta)$ for each $\sigma\in{\Gamma_{AC}(\eta)}$. The linearity of $\widetilde{f}$ is clear.
Without loss of generality, we can choose a partition $P=\{t_{0},...,t_{n}\}$ of $[a,b]$ such that there exist families of charts $\{(\varphi_{i},U_{i}):i\in\{1,...,n\}\}$ and $\{(\phi_{i},V_{i}):i\in\{1,...,n\}\}$ that verify the definition of absolute continuity for $\eta$ and $f\circ\eta$ respectively, such that
@@
For $\sigma\in{\Gamma_{AC}(\eta)}$ and for each $i\in\{1,...,n\}$ we denote $\eta_{i}=\eta|_{[t_{i-1},t_{i}]}$ and $\sigma_{i}=\sigma|_{[t_{i-1},t_{i}]}$.
Moreover, for each $i\in\{1,...,n\}$ we define the maps
$F_{i}:AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\tau\mapsto\left(d\phi_{i}\circ Tf\circ T\varphi_{i}^{-1}\right)\circ(\varphi_{i}\circ\eta_{i},\tau)$
@@
$F:\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{M})\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{N}),\quad\left(\tau_{i}\right)_{i=1}^{n}\mapsto\Big{(}F_{i}(\tau_{i})\Big{)}_{i=1}^{n}$
@@
Let $i\in\{1,..,n-1\}$. We denote
$\tau_{i}:$ $=F_{i}(d\varphi_{i}\circ\sigma_{i})$
$=(d\phi_{i}\circ Tf\circ d\varphi_{i}^{-1})(\varphi_{i}\circ\eta_{i},d\varphi_{i}\circ\sigma_{i})$
$=d\phi_{i}\circ Tf\circ\sigma_{i}$
@@
$d\phi_{i}\circ(T\phi_{i+1})^{-1}\Big{(}\phi_{i+1}\circ(f\circ\eta)(t_{i}),\tau_{i+1}(t_{i})\Big{)}$ $=d\phi_{i}\circ(T\phi_{i+1})^{-1}\Big{(}\phi_{i+1}\circ(f\circ\eta)(t_{i}),d\phi_{i+1}\circ Tf\circ\sigma_{i+1}(t_{i})\Big{)}$
$=d\phi_{i}\circ Tf\circ\sigma_{i+1}(t_{i})$
$=d\phi_{i}\circ Tf\circ\sigma_{i}(t_{i})$
@@
The topology of ${\Gamma_{AC}(\eta)}$ does not depend on the partition or charts chosen. Indeed, since the identity map $\text{id}_{M}:M\to M$ is smooth, by the previous proposition the map
@@
is smooth regardless of the partition or charts chosen.
@@
is continuous [33, Lemma 3.2], by Proposition 2.27 the inclusion map $J_{\Gamma}:\Gamma_{AC}(\gamma)\to\Gamma_{C}(\gamma)$ is also continuous. This implies that the set
@@
Let $N_{1}$ and $N_{2}$ be smooth manifolds modeled on sequentially complete locally convex spaces, $p\in[1,\infty]$ and $\operatorname{pr}_{i}:N_{1}\times N_{2}\to N_{i}$ be the $i$-projection for $i\in\{1,2\}$. If $\eta_{1}\in AC_{L^{p}}([a,b],N_{1})$ and $\eta_{2}\in AC_{L^{p}}([a,b],N_{2})$, then the map
$\mathcal{P}:\Gamma_{AC}(\eta_{1},\eta_{2})\to\Gamma_{AC}(\eta_{1})\times\Gamma_{AC}(\eta_{1}),\quad\sigma\mapsto(T\operatorname{pr}_{1},T\operatorname{pr}_{2})(\sigma)$
@@
By Proposition 2.30 the map $\mathcal{P}$ is continuous and clearly linear. Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$; let $\{(\phi_{1,i},U_{1,i})\colon i\in\{1,\ldots,n\}\}$ and $\{(\phi_{2,i},U_{2,i})\colon i\in\{1,\ldots,n\}\}$ be families of charts of $N_{1}$ and $N_{2}$, respectively, that verify the definition of absolute continuity for $\eta_{1}$ and $\eta_{2}$, respectively. Then $\eta:=(\eta_{1},\eta_{2})\colon[a,b]\to N_{1}\times N_{2}$ is $L^{p}$-absolutely continuous and it is clear that the charts $\{(\phi_{1,i}\times\phi_{2,i},U_{1,i}\times U_{2,i})\colon i\in\{1,\ldots,n\}\}$ satisfy the condition of absolute continuity for $\eta$. For $j\in\{1,2\}$, consider the linear topological embedding
$\Phi_{\eta_{j},P}\colon\Gamma_{\eta_{j}}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{j}),\;\;\tau\mapsto(d\phi_{j,i}\circ\tau|_{[t_{i-1},t_{i}]})_{i=1}^{n}$
where $E_{j}$ is the modeling space of $N_{j}$. Also
$\Phi_{\eta,P}\colon\Gamma_{\eta}\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{1}\times E_{2}),\;\;\tau\mapsto((d\phi_{1,i}\times d\phi_{2,i})\circ\tau|_{[t_{i-1},t_{i}]})_{i=1}^{n}$
is a linear topological embedding; here $T_{(x_{1},x_{2})}(N_{1}\times N_{2})$ is identified with $T_{x_{1}}N_{1}\times T_{x_{2}}N_{2}$. For $i\in\{1,\ldots,n\}$, let
$\alpha_{i}\colon AC_{L^{p}}([t_{i-1},t_{i}],E_{1})\times AC_{L^{p}}([t_{i-1},t_{i}],E_{2})\to AC_{L^{p}}([t_{i-1},t_{i}],E_{1}\times E_{2})$
be the map taking a pair $(f_{1},f_{2})$ of functions to the function $t\mapsto(f_{1}(t),f_{2}(t))$; we know that $\alpha_{i}$ is an isomorphism of topological vector spaces. Then also
$\alpha\colon\left(\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{1})\right)\times\left(\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{2})\right)\to\prod_{i=1}^{n}AC_{L^{p}}([t_{i-1},t_{i}],E_{1}\times E_{2}),$
$((f_{i})_{i=1}^{n},(g_{i})_{i=1}^{n})\mapsto(\alpha(f_{i},g_{i}))_{i=1}^{n}$
is an isomorphism of topological vector spaces. If $(f_{i})_{i=1}^{n}$ is in the image of $\Phi_{\eta_{1},P}$ and $(g_{i})_{i=1}^{n}$ is in the image of $\Phi_{\eta_{2},P}$, then $\alpha((f_{i})_{i=1}^{n},(g_{i})_{i=1}^{n})$ is in the image of $\Phi_{\eta,P}$, as the compatibility at the endpoints can be checked by considering the components in $E_{1}$ and $E_{2}$. We can therefore define a function
$\Theta:=\Phi_{\eta,P}^{-1}\circ\alpha\circ(\Phi_{\eta_{1},P}\times\Phi_{\eta_{2},P})\colon\Gamma_{\eta_{1}}\times\Gamma_{\eta_{2}}\to\Gamma_{\eta},$
which is continuous and linear. We readily check that $\mathcal{P}(\Theta(\sigma,\tau))=(\sigma,\tau)$ for all $\sigma\in\Gamma_{\eta_{1}}$ and $\tau\in\Gamma_{\eta_{2}}$. Hence $\mathcal{P}$ is surjective and thus bijective, with $\mathcal{P}^{-1}=\Theta$ a continuous map. ∎
@@
$\rho:{\Gamma_{AC}(\eta)}\to\prod_{i=1}^{n}\Gamma_{AC}\left(\eta|_{[t_{i-1},t_{i}]}\right),\quad\sigma\mapsto\left(\sigma|_{[t_{i-1},t_{i}]}\right)_{i=1}^{n}$
@@
Let $j\in\{1,...,n\}$, If $P_{j}=\{t_{j-1},t_{j}\}$, since $\eta_{j}:=\eta|_{[t_{j-1},t_{j}]}\in AC_{L^{p}}([t_{j-1},t_{j}],N)$, we have
$\rho=\left(\prod_{i=1}^{n}\Phi_{\eta_{i},P_{i}}^{-1}\right)\circ\Phi_{\eta,P}.$
@@
$\left(\rho|^{\text{Im}(\rho)}\right)^{-1}=\Phi_{\eta,P}^{-1}\circ\left(\prod_{i=1}^{n}\Phi_{\eta_{i},P_{i}}\right).$
@@
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. If $g:[c,d]\to[a,b]$ is the map as in Lemma 2.15, then $\eta\circ g\in AC_{L^{p}}([c,d],N)$ for each $\eta\in AC_{L^{p}}([a,b],N)$. Moreover, if $\eta\in AC_{L^{p}}([a,b],N)$, then the map
@@
Let $\eta\in AC_{L^{p}}([a,b],N)$. We will show first that $\eta\circ g\in AC_{L^{p}}([c,d],N)$. Let $P=\{t_{0},...,t_{n}\}$ be a partition of $[a,b]$ and $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ charts of $N$ that verify the definition of absolute continuity for $\eta$. Since $g$ is a strictly increasing function, we can define a partition $Q=\{s_{1},...,s_{n}\}$ of $[c,d]$ such that $g(s_{i})=t_{i}$, for each $i\in\{1,...,n\}$. Moreover, given that
@@
which are continuous by Lemma 2.15. Considering the topological embeddings $\Phi_{\eta,P}$ and $\Phi_{\eta\circ g,Q}$ (as in Proposition 2.27) with the same family of charts, if $(\tau_{i})_{i=1}^{n}\in\text{Im}(\Phi_{\eta,P})$ we have
@@
for each $i\in\{1,...,n-1\}$. If $G=(G_{1}\times...\times G_{n})$, then $\left(G\circ\Phi_{\eta,P}\right)(\sigma)\in\text{Im}(\Phi_{\eta\circ g,Q})$ and
@@
Hence $L_{g}$ is continuous and clearly linear. ∎
@@
is continuous and linear in the first argument. Moreover, for each $t\in[a,b]$, the point evaluation map
@@
is continuous and the evaluation map $\tilde{\epsilon}_{t}:\Gamma_{C}(\eta)\to TN$, $\sigma\mapsto\sigma(t)$ is smooth for each $t\in[a,b]$ (see [3]). Then $\epsilon=\tilde{\epsilon}\circ(J_{\Gamma}\times\text{Id}_{\mathbb{R}})$ and $\epsilon_{t}=\tilde{\epsilon}_{t}\circ J_{\Gamma}$, where $J_{\Gamma}:{\Gamma_{AC}(\eta)}\to\Gamma_{C}(\eta)$ is the inclusion map, which is continuous linear by Remark 2.32. ∎
@@ 3 Manifolds of absolutely continuous functions
Every paracompact finite-dimensional smooth manifold admits a local addition. If a smooth manifold $N$ admits a local addition, then also its tangent manifold $TN$ admits a local addition [3, Lemma A.11]. Moreover, each manifold which admits a local addition also admits a normalized local addition [3, Lemma A.14]. From now we will assume that each local addition is normalized.
@@
$\mathcal{V}_{\eta}:=\{\sigma\in{\Gamma_{AC}(\eta)}:\sigma([a,b])\subseteq\Omega\}$
@@
The following lemma is just an application of [4, Lemma 10.1] to our particular case.
@@
Let $E$ and $F$ sequentially complete locally convex spaces, $U\subseteq E$ open and $f:U\to F$ a map. If $F_{0}\subseteq F$ is a closed vector subspace and $f(U)\subseteq F_{0}$, then $f:U\to F$ is smooth if and only if $f|^{F_{0}}:U\to F_{0}$ is smooth.
@@
We endow $AC_{L^{p}}([a,b],N)$ with the final topology with respect to the family of maps
$\Psi_{\eta}\colon{\mathcal{V}}_{\eta}\to{\mathcal{U}}_{\eta},\quad\text{for each }\eta\in AC_{L^{p}}([a,b],N).$
If we define the maps $\Psi_{\eta}^{C}\colon{\mathcal{V}}_{\eta}^{C}\to{\mathcal{U}}_{\eta}^{C}$ on the space of continuous functions $C([a,b],N)$ for each $\eta\in C([a,b],N)$ as in Remark 3.3, then the final topology on $C([a,b],N)$ with respect to this maps coincides with its compact-open topology. Hence the inclusion map
@@
$\mathcal{U}_{\eta}^{C}:=\{\gamma\in C([a,b],N):(\eta,\gamma)([a,b])\subseteq\Omega^{\prime}\}$
is open in $C([a,b],N)$, whence